OpenStudy (anonymous):

[Multiple Integration] Integration of "xcos(xy)" with respect to dydx. The bounds are 0->1 for x. And 0->pi/2 for y. I tried doing this, but I don't think I have the first integration right. From my understanding, the first "x" stays as is since I am integrating w.r.t dy first (constant). Cosine changes to sine. But integrating the inside (xy) part confuses me. Am I supposed to use u-sub? In the second integral I think I have to use Itegration by Parts. I'm a bit rusty. Can someone help me?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (amistre64):

\[\int\int xcos(xy)dydx\] \[\int\left(\int xcos(xy)dy\right)dx\] \[\int[\sin(xb)-\sin(xa)]dx\] etc

OpenStudy (amistre64):

sin(3y) derives to 3cos(3y) ... so assume x is a constant wrty